Ponderosa pines grow about 1 1/2 feet each year. If a pine is now 17 feet tall, about how long will it take the tree to become 33 1/2 feet tall?
Answer by stanbon(75887)

Ponderosa pines grow about 1 1/2 feet each year. If a pine is now 17 feet tall, about how long will it take the tree to become 33 1/2 feet tall?
---------------
h(t) = 17 + (3/2)t
-----
33 1/2 = 17 + (3/2)t
----
16 1/2 = (3/2)t
----
(33/2) = (3/2)t
t = 11 years
=====================
Cheers,
Stan H.
=====================